OpenStudy (anonymous):

Find a vector parametric equation $$\vec{r}(t)$$ for the line through the points P=(−1,−4,2) and Q=(−1,1,−3) for the given conditions on the parameter t. If $$\vec{r}(3)=P$$ and $$\vec{r}(6)=Q$$

OpenStudy (anonymous):

I tried some guess and check but cant seem to get it

OpenStudy (anonymous):

is there an algebraic way to do this im not getting?

ganeshie8 (ganeshie8):

r(3) = P useually you get r(0) = P so this is like you're shifting the line to right by 3 units

ganeshie8 (ganeshie8):

whats the transformation rule for shifting the graph of f(t) to right by 3 units ?

ganeshie8 (ganeshie8):

f(t) --> f(t-3) yes ?

OpenStudy (anonymous):

hmm im not aware of that rule

ganeshie8 (ganeshie8):

Okay, side question : consider f(x) = x^2 whats the new equation after shifting above graph to right by 3 units ?

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

i guess (x-3)^2?

ganeshie8 (ganeshie8):

see if you can go thru this quick http://www.mathsisfun.com/sets/function-transformations.html

OpenStudy (anonymous):

f(x) = (x-3)^2

ganeshie8 (ganeshie8):

Yes! easy, isnt it ?

OpenStudy (anonymous):

yes, I learned that I think a long time ago and forgot

ganeshie8 (ganeshie8):

happens :) see if below satisfies the given conditions now : \[\langle -1, -4, 2\rangle + \frac{t-3}{3}\langle 0, 5, -5\rangle \]

OpenStudy (anonymous):

it does, but why divide by 3?

ganeshie8 (ganeshie8):

to slow down the speed by a factor of 3 because you want to take 3 units in going from P to Q : r(3) = P r(6) = Q 6-3 = 3

ganeshie8 (ganeshie8):

we wont be dividing by 3 if the conditions were like below : r(3) = P r(4) = Q

OpenStudy (anonymous):

ahhh, interesting

OpenStudy (anonymous):

tyvm :), I have a part 3 to this question which im not even sure what its asking..

OpenStudy (anonymous):

If the points P and Q correspond to the parameter values t=0 and t=−4, respectively

OpenStudy (anonymous):

i will post up a new question

OpenStudy (anonymous):

I mean I think I understand each question by itself, but I would think it should help me do the next question, but I get stuck on every one of them

ganeshie8 (ganeshie8):

its okay, lets try and figure out here :)

OpenStudy (anonymous):

I mean I understand each question after getting help

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so I think it means that if i have r(0) it will give me P and if I have r(4) it will give me Q?

ganeshie8 (ganeshie8):

`If the points P and Q correspond to the parameter values t=0 and t=-4, respectively` this is the question ?

OpenStudy (anonymous):

yes, thats t = -4 sorry

ganeshie8 (ganeshie8):

so you want to reach point P at t=0 and reach point Q at t = -4

ganeshie8 (ganeshie8):

think of "t" as time if it helps..

ganeshie8 (ganeshie8):

clearly you're going back in time

ganeshie8 (ganeshie8):

try this : \[\langle -1, -4, 2\rangle + \frac{-t}{4}\langle 0, 5, -5\rangle \]

OpenStudy (anonymous):

yep that did work, wow, why cant i think of this :(

OpenStudy (anonymous):

i did try t/4

ganeshie8 (ganeshie8):

you just need to see it worked one time, im sure you will figure it out by yourself next time :)

OpenStudy (anonymous):

i hope so :), ty so much, really appreciate it

ganeshie8 (ganeshie8):

yw!